Fermi National Laboratory


Accelerator Update for November 23, 2005 - November 28, 2005

Wednesday November 23

The day shift began with the Tevatron (TeV) cryo system warming up, with Operations monitoring the Antiproton Source (Pbar) stacking, and with MiniBooNE, NuMI, and the Meson MTest and MCenter experiments taking beam.

The Main Injector (MI) and Pbar conducted studies and technicians and experts began work on the TeV sector B1 vacuum and cryo system work. By the middle of the day Pbar had resumed stacking and the Recycler stashed antiprotons. Machine experts continued with studies. Interlock experts conducted safety system tests in Meson.

Thursday November 24

The midnight shift began with the TeV cryo warm up in progress, with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

Experts tuned the Booster, MI, and Switchyard beam. The Recycler stashed and studies continued. A Booster kicker (MKS08) failed and was repaired.

Friday November 25

The midnight shift began with TeV sector B1 cryo warming up, with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

The B17 cryo warm up was completed at 4:30 AM. Cryo techs began preparing B1 cryo system for the magnet repair and spool piece replacement early in the morning. Operators traced a problem with stacking to the lithium lens. An expert reported that the problem was in the tunnel. Pbar conducted reverse proton studies and MI conducted 2.5 Mhz studies. The Recycler had stochastic cooling problems. Operators were investigating H- Source problems at the end of the evening shift.

Saturday November 26

The midnight shift began with TeV B17 work, with Operations monitoring the stash, with Pbar and MI conducting studies, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

The Recycler stashed antiprotons and then conducted cooling studies. An expert reported that the brushless generator for the H- Source had failed. Vacuum techs continued with the B17 bypass and leak check. Booster experts replaced a Thyratron tube for a kicker (MKS07).

Sunday November 27

The midnight shift began with TeV B17 work, with Operations monitoring a small stash, with Pbar, Recycler, and MI conducting studies, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

Vacuum techs reported that the TeV B17 pumpdown was successful. By the end of the evening shift, the techs were preparing to run a video-scope into the bypass.

Monday November 28

The midnight shift began with cryo experts scoping the B17 bypass, with Pbar, Recycler, and MI conducting studies, with MiniBooNE, NuMI, MTest, and MCenter taking beam.

Experts reported that the B17 bypass was okay. A Booster RF station (BRF16) tripped and wouldn't reset. A Linac quadrupole power supply (QPS107) tripped off and wouldn't reset. Linac techs had the QPS107 back on by the end of the midnight shift.

The Plans for Monday

The plans for today are to continue with the B17 work. The experts estimate that the dipole magnet and spool piece replacement will be completed and the cryo system cooled down by Wednesday, December 7. The Run Coordinator is considering having a 12-hour downtime on Tuesday for maintenance and repair, but this hasn't been confirmed.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/28/2005   email Fermilab